OpenStudy (jimbo69):

The state of Wyoming has roughly the shape of a rectangle that is 3.6 x 102 miles long and 2.8 x 102 miles high. What is the approximate area of Wyoming? Hint: The area of a rectangle is product of its length and width.

OpenStudy (studygurl14):

Multiply \(3.6\times 10^2\) and \(2.8\times 10^2\) together. Remember that you can rearrange it to \((3.6\times 2.8)(10^2 \times 10^2)\)